# Revealing The Secret Power: How Algorithms Can Influence Content Visibility on Social Media

### v2

@@ 5.1 RQ1: Content Level
5.2 RQ2: User Level
Inferring Users’ Ideology. To perform account-level analysis, we first need to infer users’ attributes, such as their stance on the debated issue. To this end, we use latent ideology estimation (see Section 4.4) to classify users’ ideological stances. Given a set of prominent accounts known as influencers (see Section 4.3), this technique allows us to infer users’ stances based on interaction information. Matrix decomposition maps users who interacted with a similar set of influencers close to each other in a real space, while users with fewer common influencers are positioned farther apart.
Due to the absence of retweet statistics for the 2024 US Elections dataset, we estimate latent ideology using retweet interactions for the Ukraine-Russia and reply interactions for the US Elections datasets. The use of replies does not affect the ability to infer users’ stances as shown by the influencers’ inferred stances in Figure 6. Both datasets exhibit two communities positioned at opposite ends of the $[-1,1]$ space, revealing polarized debates with bimodal opinion distributions, as also highlighted by Baqir et al. [4] in the context of the Ukraine-Russia war.
In each debate, we categorize accounts into two stances: supporters (ideology score $<0$) or opponents of military aid (ideology score $>0$), and supporters (ideology score $>0$) or opponents of Donald Trump (ideology score $<0$). We then calculate the average p-score for users and influencers in our dataset. To ensure the robustness of our results, we only consider accounts with more than five original tweets.
@@ 5.2 RQ2: User Level
Visibility vs Ideology. After inferring user opinions, we analyze whether the visibility of the account is influenced by the ideology they support. Figure 7 illustrates the distributions of the average p-score for users and influencers representing opposing ideological stances in the two debates. Influencers tend to have higher average p-scores than users in both debates, likely due to their prominent role and greater popularity. However, there is little difference between the p-score distributions of users and influencers when comparing ideological stances in both the Ukraine-Russia and US Elections datasets. For the Ukraine-Russia dataset, the median p-scores are 0.118 for supporters vs. 0.136 for opponents among users, and 0.479 for supporters vs. 0.499 for opponents among influencers. For the US Elections dataset, the median p-scores are 0.049 for Trump supporters vs. 0.048 for opponents among users, and 0.280 for Trump supporters vs. 0.358 for opponents among influencers. These results are confirmed by the Mann-Whitney U test, which shows significantly higher performance for influencers compared to users (p-value $<$ 0.0005 for both supporter and opponent influencers versus users; p-value $<1.30\times 10^{-7}$ for pro- and anti-Trump influencers versus users). The test also reveals comparable visibility between users and influencers of different stances (p-value $>$ 0.07 for the Ukraine-Russia case and p-value $>$ 0.29 for the US elections).
@@
Influences p-score distributions. Although we have previously ruled out the penalization of specific ideologies, visibility alterations may still occur for specific accounts. This is suggested by the presence of outliers among influencers in Figure 7, which may indicate accounts experiencing either reduced or increased content circulation due to algorithmic visibility adjustments. To investigate the presence of algorithmic interventions and to mitigate the influence of potential confounding factors, we exclusively focus on influencers and divide them based on their popularity in different tiers, as done in recent work [37]. Figure 6 displays the p-score distribution for the accounts used as influencers in the analysis of the two debates that posted more than five original pieces of content, with accounts grouped according to the popularity tier they belong to.
@@
For both datasets, within the same group, influencers tend to exhibit distributions with similar shapes but different medians, reinforcing the possibility of algorithmic intervention at the individual level. We also observe that some influencers exhibit multiple peaks in their p-score distributions. This pattern could be attributed to the use of different types of content, each achieving varying levels of visibility, as illustrated in Figure 1. Alternatively, it may reflect temporal fluctuations in visibility due to adjustments in the suggestion algorithm. Notably, in the Ukraine-Russia war debate, mfa_russia and RT_com —- accounts associated with the Ministry of Foreign Affairs and a Russian state-sponsored news outlet – significantly underperform relative to other accounts within the same tier. In the US Elections debate, the accounts linked to Joe Biden and Kamala Harris exhibit p-score distributions skewed lower than that of Donald Trump.
One might argue that the circulation of content also depends on how much other users engage with it; thus, a more active user base could enhance the diffusion of the author’s content. To assess this, Figure 8 presents a 2-D density plot of the p-score versus the ratio of retweets to views. This ratio serves as a proxy for quantifying how actively consumers re-share content after being exposed to it. In particular, Figure 8(a) reveals that RT_com and KyivIndependent have rather similar user bases in terms of active sharing, as highlighted by the top marginal distribution (Mann-Whitney U test p-value: 0.093). However, they display significant differences in their p-score distributions (Mann-Whitney U test p-value: $<2.2\times 10^{-16}$, median p-score: KyivIndependent=0.084, RT_com=0.007). Given the similarities across various dimensions, these findings strongly support the hypothesis that content from RT_com has been subject to visibility reduction.
In the case of the US Elections, the separation between Trump-Harris distributions is less clear (Figure 8(b)), as the two 2D densities overlap in some areas. Nevertheless, their marginal distributions differ significantly along both dimensions. Specifically, the distribution of retweet count over view count reveals that Donald Trump’s audience is less prone to explicitly endorse his content through active engagement compared to Kamala Harris’s audience (median retweets/views: 0.0025 for Trump, 0.0043 for Harris, Trump vs Harris Mann-Whitney U test p-value=1, Harris vs Trump Mann-Whitney U test p-value $=4.543\times 10^{-15}$). However, the p-score marginal distributions show that most of the content produced by Donald Trump achieves greater visibility than that produced by Kamala Harris (p-score median: 0.13 Trump, 0.03 Harris, Trump vs Harris Mann-Whitney U test p-value $<2.2\times 10^{-16}$). Hence, despite having a less active user base, Donald Trump’s content experienced increased visibility, suggesting the possible influence of algorithmic intervention.
@@
Figure 8: P-score versus retweets per impression count for content produced by two news agencies with opposing ideological stances.
@@
5.3 RQ3: Network Level
@@
The analysis indicates an absence of correlation between individual p-scores and the average p-scores of retweeters across both account classes and ideological stances for both datasets. Thus, users who interacted with accounts with reduced visibility do not necessarily experience decreased visibility themselves, regardless of account type, stance, or debate. This is further supported by Pearson’s correlation test, which yielded a maximum correlation value of 0.13 for the Influencer-Supporter case in the Ukraine-Russia dataset and a value of 0.094 in the US elections debate.
6 Discussion
The visibility of content online is heavily shaped by recommendation algorithms, which determine what users see based on various engagement metrics and platform-specific goals [23]. Many platforms report reducing visibility as part of their moderation strategies to limit the spread of harmful or borderline content, purportedly to safeguard public discourse [13]. However, the lack of transparency around these interventions and the policies driving them contributes to a growing distrust towards platforms [33]. On one hand, interventions can effectively reduce the spread of harmful content or limit the influence of problematic users, including bots or coordinated disinformation campaigns. On the other hand, there is no guarantee that the same mechanisms are not used to silence minority voices or suppress narratives contrary to corporate interests. These interventions are subtle, far less noticeable than bans or content removal, and users may remain unaware that the reach of their posts is being reduced [13].
The role of algorithms. Our work investigated visibility reduction in the context of two Twitter datasets related to two polarized topics (the Ukraine-Russia war and the 2024 US Elections) using the content views metric. This allowed us to directly observe the effects of algorithmic interventions and assess how visibility reductions vary based on several factors. In doing so, we analyze both the behavior of recommendation algorithms at a granular level and how views can help identify otherwise unnoticed visibility suppression.
Data availability. Practices like shadow banning can significantly influence online debates and shape public opinion. As a result, the ability to study these interventions is crucial. However, while data availability is essential in doing so, in recent years, access to social media data for research purposes has become increasingly restricted [24], exacerbating the opacity of platforms and further eroding public trust. For instance, there is growing concern that social media companies could reduce content visibility to reduce the spread of misinformation but may be choosing not to do so [1].
@@ 6 Discussion
Longitudinal analysis. The comparison between multiple datasets collected during different time periods and on different topics offers also helped us highlight the evolution of visibility depending on the context considered. For example, although the algorithm tends to reduce the visibility of tweets containing URLs, the strength of such interventions appears to differ between the two cases. Similarly, the differences observed in the comparative analysis of visibility at the individual level may result from variations in the behavior of the recommendation algorithm.
@@
Additional factors. Finally, other factors influencing content visibility, including coordinated behaviors, user engagement strategies, and the topic under discussion, should also be taken into account. While these factors can exploit algorithmic design to boost content visibility, our methodology proves effective in detecting visibility disparities between different types of content and user behavior independently from the origin of these differences.
7 Conclusion
@@
This paper examined differences in content visibility on Twitter/X – at content, user, and network levels – in the context of the Ukraine-Russia war and the 2024 US Presidential Elections We found evidence that Twitter/X consistently favors tweets that do not contain links to external websites. While the platform does not systematically alter content visibility based on political views, the factuality of the referenced sources, or ideological stances, we do find evidence of visibility alterations at the individual account level, as illustrated by the cases of RT.com versus The Kyiv Independent and Donald Trump versus Kamala Harris. Our work demonstrates how visibility alterations can be detected using engagement metrics like view counts, an underexplored and often unavailable metric.
@@
Ethics Considerations. Our work relies on publicly available data obtained from other studies [4, 3], retrieved using the Twitter public API or by scraping public data, and augmented with publicly available tools. Therefore, it is not considered human subject research by our institutions. To preserve anonymity, we do not report any information about the users in the dataset other than those of public entities like news organizations. For accounts that play a prominent role in the debate (i.e., influencers), we anonymize those with fewer than 50,000 followers, aligning with previous research [11].

### v3

@@ V-B RQ2: User Level
Visibility vs Ideology. After inferring user opinions, we analyze whether the visibility of the account is influenced by the ideology they support. Figure 8 illustrates the distributions of the average p-score for users and influencers representing opposing ideological stances in the two debates. Influencers tend to have higher average p-scores than users in both debates, likely due to their prominent role and greater popularity. However, there is little difference between the p-score distributions of users and influencers when comparing ideological stances in both the Ukraine-Russia and US Elections datasets. For the Ukraine-Russia dataset, the median p-scores are 0.118 for supporters vs. 0.136 for opponents among users, and 0.479 for supporters vs. 0.499 for opponents among influencers. For the US Elections dataset, the median p-scores are 0.049 for Trump supporters vs. 0.048 for opponents among users, and 0.280 for Trump supporters vs. 0.358 for opponents among influencers. These results are confirmed by the Mann-Whitney U test, which shows significantly higher performance for influencers compared to users (p-value << 0.0005 for both supporter and opponent influencers versus users; p-value <1.30e07<1.30e-07 for pro- and anti-Trump influencers versus users). The test also reveals comparable visibility between users and influencers of different stances (p-value >> 0.07 for the Ukraine-Russia case and p-value >> 0.29 for the US elections).
@@
Influences p-score distributions. Although we have ruled out the penalization of specific ideologies, visibility alterations may still occur for specific accounts. This is suggested by the presence of outliers among influencers in Figure 8, which may indicate accounts experiencing either reduced or increased content circulation due to algorithmic visibility adjustments. To investigate the presence of algorithmic interventions and to mitigate the influence of potential confounding factors, we exclusively focus on influencers and divide them based on their popularity in different tiers, as done in recent work [40]. Figure 7 displays the p-score distribution for the accounts used as influencers in the analysis of the two debates that posted more than five original pieces of content, with accounts grouped according to the popularity tier they belong to.
@@
For both datasets, within the same group, influencers tend to exhibit distributions with similar shapes but different medians, reinforcing the possibility of algorithmic intervention at the individual level. We also observe that some influencers exhibit multiple peaks in their p-score distributions. This pattern could be attributed to the use of different types of content, each achieving varying levels of visibility, as illustrated in Figure 1. Alternatively, it may reflect temporal fluctuations in visibility due to adjustments in the suggestion algorithm. Notably, in the Ukraine-Russia war debate, mfa_russia and RT_com (accounts associated with the Ministry of Foreign Affairs and a Russian state-sponsored news outlet, respectively) significantly underperform relative to other accounts within the same tier. In the US Elections debate, the accounts linked to Joe Biden and Kamala Harris exhibit p-score distributions skewed lower than that of Donald Trump.
@@
Figure 9: P-score versus retweets per impression count for content produced by two news agencies with opposing ideological stances in the Ukraine-Russia war debate (a) and by the two presidential candidates in the US Election discussion (b). The volume of content produced by the four accounts are 553(KyivIndependent), 258(RT_com), 223(Donald J. Trump) and (467)Kamala Harris.
@@
One might argue that the circulation of content also depends on how much other users engage with it; thus, a more active user base could enhance the diffusion of the author’s content. To assess this, Figure 9 presents a 2-D density plot of the p-score versus the ratio of retweets to views. This ratio serves as a proxy for quantifying how actively consumers re-share content after being exposed to it. In particular, Figure 9(a) reveals that RT_com and KyivIndependent have rather similar user bases in terms of active sharing, as highlighted by the top marginal distribution (Mann-Whitney U test p-value: 0.093). However, they display significant differences in their p-score distributions (Mann-Whitney U test p-value: <2.2e16<2.2e-16, median p-score: KyivIndependent=0.084, RT_com=0.007). Given the similarities across various dimensions, these findings strongly support the hypothesis that content from RT_com has been subject to visibility reduction.
In the case of the US Elections, the separation between Trump-Harris distributions is less clear (Figure 9(b)), as the two 2D densities overlap in some areas. Nevertheless, their marginal distributions differ significantly along both dimensions. Specifically, the distribution of retweet count over view count reveals that Donald Trump’s audience is less prone to explicitly endorse his content through active engagement compared to Kamala Harris’s audience (median retweets/views: 0.0025 for Trump, 0.0043 for Harris, Trump vs Harris Mann-Whitney U test p-value=1, Harris vs Trump Mann-Whitney U test p-value =4.543e15=4.543e-15). However, the p-score marginal distributions show that most of the content produced by Donald Trump achieves greater visibility than that produced by Kamala Harris (p-score median: 0.13 Trump, 0.03 Harris, Trump vs Harris Mann-Whitney U test p-value <2.2e16<2.2e-16). Hence, despite having a less active user base, Donald Trump’s content experienced increased visibility, suggesting the possible influence of algorithmic intervention.
V-C RQ3: Network Level
@@
The analysis indicates an absence of correlation between individual p-scores and the average p-scores of retweeters/repliers across both account classes and ideological stances for both datasets. Thus, users who interacted with accounts with reduced visibility do not necessarily experience decreased visibility themselves, regardless of account type, stance, or debate. This is further supported by Pearson’s correlation test, which yielded a maximum correlation value of 0.13 for the Influencer-Supporter case in the Ukraine-Russia dataset and a value of 0.094 in the US elections debate.
VI Discussion
The visibility of content online is inherently shaped by recommendation algorithms, as these determine what users see based on various engagement metrics and platform-specific goals [26]. Many platforms report reducing visibility as part of their moderation strategies to limit the spread of harmful or borderline content, purportedly to safeguard public discourse [16]. However, algorithmic visibility reduction may occur for several reasons, for instance, to penalize low-performing content or to prioritize user retention.
Regardless, the lack of transparency around these interventions contributes to a growing distrust towards platforms [36]. Moreover, while interventions can reduce the spread of harmful content or limit the influence of problematic users (including bots or coordinated disinformation campaigns), there is no guarantee that the same mechanisms are not used to silence minority voices or suppress narratives contrary to corporate interests. These interventions are subtle, far less noticeable than bans or content removal, and users may remain unaware that the reach of their posts is being reduced [16].
Our work investigated visibility reduction in the context of two Twitter datasets related to two polarized topics (the Ukraine-Russia war and the 2024 US Elections) using the content views metric. This allowed us to directly observe the effects of algorithmic interventions and assess how visibility reductions vary based on several factors. In doing so, we analyze both the behavior of recommendation algorithms at a granular level and how views can help identify otherwise unnoticed visibility suppression.
Data availability. Practices like shadow banning can significantly influence online debates and shape public opinion. As a result, the ability to study these interventions is crucial. However, while data availability is essential in doing so, in recent years, access to social media data for research purposes has become increasingly restricted [27], exacerbating the opacity of platforms and further eroding public trust. For instance, there is growing concern that social media companies could reduce content visibility to reduce the spread of misinformation, but may be choosing not to do so [1].
@@ VI Discussion
Longitudinal analysis. The comparison between multiple datasets collected during different periods and on different topics also helped us highlight the evolution of visibility depending on the context considered. For example, although the algorithm tends to reduce the visibility of tweets containing URLs, the strength of such interventions appears to differ between the two cases. Similarly, the differences in the visibility analysis at the individual level may result from variations in the behavior of the recommendation algorithm.
@@
Additional factors. Finally, other factors influencing content visibility, including coordinated behaviors, user engagement strategies, and the topic under discussion, should also be taken into consideration. While these factors can exploit algorithmic design to boost visibility, our methodology proves effective in detecting visibility disparities between types of content and user behavior independently from the origin of these differences. In other words, alternative explanations for our findings are, arguably, far less likely.
For instance, one possible reason for visibility alteration could be related to bots, however, our visibility metric adjusts for follower count to account for potential bot-driven inflation. Furthermore, our network analysis is designed to detect clusters of users with altered visibility, including those that may be the result of coordinated behavior. Since no penalized communities were identified, such groups, if they exist, either evade detection or do not benefit from increased visibility, making coordinated manipulation an unlikely explanation. At the user level, we include retweets to capture engagement-driven visibility; nevertheless, users with similar roles and activity levels still exhibit significant differences in visibility, which suggests that user-level factors alone cannot explain the observed penalization. Lastly, the systematic reduction in visibility of content containing URLs in both debates, regardless of the type of resource referenced, strongly suggests algorithmic intervention rather than varying user interest.
Ethics Considerations. Our work relies on publicly available data obtained from other studies [5, 4], retrieved using the Twitter public API or by scraping public data, and augmented with publicly available tools. As such, it is not considered human subject research by our institutions. We acknowledge that one of our datasets was collected by [4] using a custom scraper, which may potentially contravene Twitter’s terms of service (ToS). However, we believe our research to be in the public interest, as studying platform manipulation arguably outweighs ToS concerns, in a manner not dissimilar to, e.g., violating robots.txt when crawling cybercrime forums [8]. In general, ethics researchers have made the case that ethics might not depend on breaking the ToS of the platforms [14].
Furthermore, to preserve anonymity, we do not report any information about the users in the dataset other than those of public entities like news organizations. For accounts that play an important role in the debate (i.e., influencers), we anonymize those with fewer than 50,000 followers, aligning with previous research [13]. Overall, we believe that our work aligns with the Menlo Report [2]’s principles of Beneficence, Respect for Persons, and Public Interest.
VII Conclusion
@@
This paper examined differences in content visibility on Twitter/X – at content, user, and network levels – in the context of the Ukraine-Russia war and the 2024 US Presidential Elections. We relied on view counts, a largely unexplored feature of Twitter/X posts, to estimate content circulation and detect forms of visibility penalization that are not captured by techniques previously used in the literature. We found evidence that Twitter/X consistently favors tweets that do not contain links to external websites. While the platform does not systematically alter content visibility based on political views, the factuality of the referenced sources, or ideological stances, we do find evidence of visibility alterations at the individual account level, as illustrated by the cases of RT.com versus The Kyiv Independent and Donald Trump versus Kamala Harris. Our work demonstrates how visibility alterations can be detected using engagement metrics like view counts, an underexplored and often unavailable metric.
As mentioned, our work mainly focuses on algorithmic penalization in a broad sense, while we do not aim to determine whether/how this may be ideologically motivated.
Recommendations. We believe that the methods proposed in this paper can help researchers and civil society alike improve transparency by uncovering penalization practices adopted by social media platforms, especially if they might end up infringing on free speech. Concretely, non-profits, research outfits, and fact-checkers can leverage our methods to raise awareness and increase transparency by publishing dashboards of content, keywords, and domains that are anomalously penalized, and record historical information about potential changes in visibility dynamics in public reports.
@@
Acknowledgments. A.G. and G.S. gratefully acknowledge support from, respectively, CY4GATE and NSF Grant CNS-2419829.